use long division to express the function x2-3x+7/x+2 as the sum of a polynomial function and a rational function of the form A/x+2, where a is constant


1
Expert's answer
2021-11-10T13:15:10-0500


so "\frac{x^2-3x+7}{x+2}=(x-5)+\frac{17}{x+2}"


"\implies f(x)=(x-5)+\frac{17}{x+2}"